# Continual Learning from Simulated Interactions via Multitask Prospective Rehearsal for Bionic Limb Behavior Modeling

### v3

@@ 1 Introduction
Lower limb amputation and neuromuscular disorders detrimentally affect natural locomotion, compromising individuals’ quality of life (Windrich et al., 2016). Those afflicted often depend on assistive technologies like prosthetics or orthoses to regain daily mobility (Windrich et al., 2016). However, conventional passive prosthetics often fall short in replicating natural gait during diverse activities, from basic transitions like standing from a seated position to more dynamic actions like running or navigating slopes (Windrich et al., 2016). The advent of powered prosthetics, equipped with integrated motors, promises a more naturalistic gait. Nonetheless, this advancement requires a model capable of effectively approximating the complexities of human gait synergy to accurately estimate the necessary motor commands. While finite-state machines are adequate for rudimentary scenarios (Lawson, 2014; Chen et al., 2015), their construction becomes infeasible for finer gait phase resolution and multiple locomotion tasks (Lawson et al., 2013). An emerging solution lies in training models on human demonstrations to intuitively infer target limb motion (Dey et al., 2019; 2020a; 2021b). This methodology not only replicates gait’s inherent fluidity but also facilitates seamless transitions across varying gaits, eliminating the need for rigid rules or heuristics. Furthermore, such learning-based gait models can be robustified by training on a large and diverse set of human demonstrations Dey et al. (2020b). To achieve personalization, these models could incorporate references that exhibit anthropomorphic features similar to those of the physically impaired subject Dey & Nair (2024).
@@
Recent studies attest to the efficacy of learning-based models in gait and prosthetic behavior estimation (Dhir et al., 2018; Zabre-Gonzalez et al., 2021; Fang et al., 2020; Dey & Schilling, 2022a). However, most of these works have focused on either a single locomotion task or jointly training multiple modes in the context of multi-locomotion scenarios. Furthermore, prior research has approached multitask learning from human demonstration as a standard supervised learning task, which assumes independence across time. This oversight can cascade predictive errors, deviating significantly from the original training distribution (Kumar et al., 2022; Ross & Bagnell, 2010). Some approaches (Ross et al., 2011; Judah et al., 2014; Kelly et al., 2019) have addressed the issue of state distribution shift by using a trained model to collect additional data and label them as the model encounters new states, refining the model iteratively. However, these studies require having access to a simulator, environment, or deployment of models in the real world for data collection to estimate the expert action for the induced state. This may not be feasible or highly costly for multiple safety-critical scenarios, e.g., human-centered robotics that might entail risks of imbalance and fall. Moreover, these approaches have largely focused on single-task settings rather than continual multitask adaptation. The multifaceted dynamics of human gait, which can shift based on varying terrains, speeds, fatigue levels, or specific locomotion tasks, underscore the necessity for a model that fluidly adapts to these evolving conditions.
@@ 2 Related Work
Prosthetic behavior models. Conventional prosthetic behavior models primarily rely on state machines that deterministically generate control commands based on finite states within a gait cycle. These models segment the continuous gait cycle into discrete states (Lawson, 2014; Chen et al., 2015; Culver et al., 2023), which fails to accurately replicate the smooth and continuous movement patterns typical of human walking. To better model the fluidity of the gait cycle, learning-based regression methods (Dey et al., 2020b; 2021a; Dhir et al., 2018; Fang et al., 2020; Zabre-Gonzalez et al., 2021; Dey & Schilling, 2022b) have been utilized that directly predict the joint motion commands from the sensor states of residual body movement. However, these models generally adopt a straightforward reactive strategy, learning the correlation between the residual states of the body and the limb’s movements. Hence, these models often oversimplify the complex nature of human motor abilities, which are inherently adaptable to various tasks, diverse environments, and evolving gait patterns.
@@
Multitask adaptive learning. A principal challenge in multitask adaptive learning is the phenomenon of catastrophic forgetting (McCloskey & Cohen, 1989; Robins, 1995; French, 1999), where newly acquired knowledge can cause the loss of previously learned information. To combat this, continual learning strategies, such as regularization-based methods, architectural strategies, and experience replay (ER)(Kirkpatrick et al., 2017; Zenke et al., 2017; Smith et al., 2023) have been developed. Regularization-based methods like Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017) and Synaptic Intelligence (SI) (Zenke et al., 2017) introduce regularization terms to guide the model towards a parameter space optimized for low error across both previous and new tasks. Architectural methods such as Progressive Neural Networks (PNNs) (Rusu et al., 2016) expand the model’s structure by introducing task-specific parameters to sustain performance across tasks. Rehearsal or replay methods, including Experience Replay (ER) (Robins, 1995; Luo & Li, 2020; Li et al., 2023) and Gradient Episodic Memory (GEM) (Lopez-Paz & Ranzato, 2017), maintain a sample of previous tasks to be revisited during new task training, reinforcing the model’s exposure to the old data distribution. Rehearsal approaches have empirically proved to be the most effective among several approaches developed (Merlin et al., 2022). However, these techniques have a retrospective nature and are meant for treating data samples independently. In the context of bionic prosthetics, where predictive models interact with a dynamic environment, anticipating the impact of predictions on subsequent data inputs is critical.
@@
Model-based reinforcement learning (RL). Model-based RL focuses on developing a representation of the environment by gathering data from interactions driven by a specific policy. This constructed model of the environment facilitates a level of planning, enabling the estimation of potential outcomes for future states (Mnih et al., 2013; Silver et al., 2016; Racanière et al., 2017; Nagabandi et al., 2018; Chua et al., 2018). Typically, these methods rely on a reward signal and online interactions with the environment. With the advent of offline reinforcement (Zhou et al., 2021; Yu et al., 2020) and model-based imitation learning (IL) (Englert et al., 2013; Kidambi et al., 2021), access to the reward signals or online interactions can be bypassed (Hu et al., 2022). Both model-based RL and IL, by internalizing environmental dynamics, enable an agent to forecast outcomes and make informed decisions. Our approach is inspired by model-based RL; however, we use it as a rehearsal strategy within a multitask continual learning setting to mitigate forgetting, adapt to evolving tasks, and limit compounding errors, all without needing access to a reward signal or online interactions.
@@ 3 Method
Multitask learning via generic and task-specific modules. A core challenge in multitask learning is to exploit commonalities between the tasks while maintaining flexibility to account for the particularities of every single task. Our approach is to compose the gait model $f$ of two functions: a shared backbone $f^{s}$ across tasks, followed by task-specific layers $f_{t}$. The overall prediction for an input sample $\{\mathbf{x}_{t,i-T}...\mathbf{x}_{t,i}\}$ for task $t$ at time $i$ is $f(\mathbf{x}_{t,i-T}...\mathbf{x}_{t,i})=f_{t}[f^{s}(\mathbf{x}_{t,i-T}...\mathbf{x}_{t,i})]$. This duality enables the model to distill common features across all tasks and concurrently adapt to the unique aspects of individual tasks. We employ temporal convolutions (Oord et al., 2016; Fang et al., 2020) to model the shared backbone due to their effectiveness in handling temporal sequences. To tailor the model to specific tasks, we implement task-specific layers using a two-layer feed-forward network, providing lightweight and efficient customization for each task.
@@
Continual learning via evolving architecture and rehearsal. In (task) incremental adaptation, the model is trained incrementally with one task at a time. Thus, the model does not require data from all the tasks during training. When a new task $t$ is encountered, a new task-specific prediction layer $f_{t}$ is created and both $f_{t}$ and $f^{s}$ are trained using task-specific data. However, this method is prone to catastrophic forgetting (French, 1999; McCloskey & Cohen, 1989), that is, the model forgets the previously trained tasks when it learns a new task. To deal with the forgetting problem, we include samples from prior tasks in the training (i.e., a ’rehearsal’ (van de Ven et al., 2020)). However, beyond prior rehearsal approaches, we do not merely copy training data from prior tasks. Instead, we augment it, incorporating prediction errors from the current model $f$ using the following scheme.
@@
Prospective model. A principal challenge in our scenario is that prediction errors at time step $k$ may negatively impact prediction performance at time step $k+1$, such that samples are not identically and independently distributed across time. In more detail, for the human demonstrations, we know that the sensor state $\mathbf{x}_{t,k}$ and desired kinematic profile $y_{t,k}$ is followed by the sensor state $\mathbf{x}_{t,k+1}$. However, if we predict a slightly different kinematic profile $\hat{y}_{t,k}=f(\mathbf{x}_{t,k-T}...\mathbf{x}_{t,k})\neq y_{t,k}$, we will also observe a different sensor state $\hat{\mathbf{x}}_{t,k+1}\neq\mathbf{x}_{t,k+1}$ in the next time step, pushing us away from the training data distribution. Such deviations can accumulate over time, which is a well-known phenomenon in imitation learning (Kumar et al., 2022; Ross & Bagnell, 2010). In general, imitation learning is a helpful metaphor for our situation, as we also wish to mimic the demonstration of human experts in a setting where independence over time does not hold. However, past theory on imitation learning has usually assumed constant bounds on the error in every time step. Such constant bounds are not realistic in our scenario with a continuous space. Unfortunately, if we slightly relax the assumption of constant bounds to Lipschitz bounds, the deviation between desired time series and predicted time series can grow exponentially, as we show in the following, simple theorem.
@@
This observation motivates the key innovation of our work, namely a prospective rehearsal. More precisely, we train a prospective model $g$ to mimic the dynamics of the system with $g(\mathbf{x}_{k},{y}_{k})\approx\mathbf{x}_{k+1}$ and imagine the subsequent inputs resulting from the predictions of $f$. We then add data tuples $(\textbf{x}^{\prime}_{k+1},{y}_{k+1})$ with $\textbf{x}^{\prime}_{k+1}=g\big{(}\mathbf{x}_{k},f(\mathbf{x}_{k-T}...\mathbf{x}_{k})\big{)}$ to our training data. These additional training data tuples take the prediction errors of $f$ into account and thus help $f$ to counteract its own prediction errors. In our scenario, ${y}_{k+1}$ is the right target for the input $\textbf{x}^{\prime}_{k+1}$ because ${y}_{k+1}$ represents a target joint profile in a prosthetic limb. This target profile should remain the same in the same gait phase, even if $\textbf{x}^{\prime}_{k+1}$ slightly deviates from $\mathbf{x}_{k+1}$. In a general scenario, other targets might be required. Theoretically speaking, our aim is to achieve a constant bound $d\big{(}g[\mathbf{x}_{k},f(\mathbf{x}_{k-T},...,\mathbf{x}_{k})],\mathbf{x}_{k+1}\big{)}\leq\epsilon$ for small $\epsilon>0$, such that classical imitation learning theory applies (Kumar et al., 2022; Ross & Bagnell, 2010) and an exponential deviation over time is avoided. We do so by minimizing $d\big{(}g[\mathbf{x}_{k},{y}_{k}],\mathbf{x}_{k+1}\big{)}$ explicitly in Eq. equation 1.
@@ 4 Experiments
Datasets and metrics. We use three real-world human gait experiments’ datasets. The first two, ENABL3S (Hu et al., 2018) and Embry (Embry et al., 2018), are publicly available from popular gait labs. ENABL3S comprises approximately 5.2 million training samples and 1.3 million test samples collected from ten subjects performing various locomotion activities, including level-ground walking, stair ascent and descent, and ambulating on an inclined walkway. From the Embry dataset, we selected data representing nine very distinct combinations of walking speeds (ranging from 0.8 m/s to 1.2 m/s) and inclines (ranging from -10 degrees to +10 degrees). This selection resulted in approximately 512,000 training samples and 104,000 test samples from ten subjects, encompassing a diverse range of movements. The third dataset is a novel collection from our lab, focusing on gait patterns of patients (transtibial amputees), recorded with a 200 Hz infrared motion capture system using twelve cameras, and includes activities like walking, stair ascent, and descent, with around 27K training and 11K testing samples. Ethical clearance for these experiments was obtained from our institutional review board. This dataset, along with the others, includes motion-related variables such as 3D angles of body joints and segments, velocities, and accelerations. To evaluate the accuracy of the desired joint profiles, $y_{k}$, we calculate the coefficient of determination ($R^{2}$), between the ground truth trajectories and model-based joint profiles.
@@
Ablations. We assess the contribution of each component on overall performance. We explore both joint and continual learning paradigms, comparing models with shared final layers against those with task-specific final layers Dey et al. (2024). Further, we consider both conventional and prospective rehearsal with each of the shared and task-specific final layer variants. In joint training, the training data is either augmented with the multitask prospective rehearsal from a validation set or simply with the original samples from the validation set itself (conventional rehearsal). In all conditions, we used the same amount of training data to maintain fairness. Augmentation with prospective rehearsal improved the performance of models with both shared and task-specific final layers. Further, the results provide evidence that task-specific final layers help to account for particularities among tasks. A comparison of performance in task-incremental learning yields similar results. The results demonstrate that the choice of a combination of prospective rehearsal and an evolving architecture, where task-specific final layers are integrated with a shared backbone, yields the best performance across the board (Table 1A–B). All following experiments are performed with task-specific final layers.
@@
Resilience to distribution shift. In some cases, we do not observe a significant improvement in the performance of the prospective rehearsal over a conventional rehearsal. To explain these results, we inspect the difference in $R^{2}$ between the prospective and conventional rehearsal versus the Jenson-Shannon (JS) distance (Endres & Schindelin, 2003) between training and test distributions (Figure 3). We observe that the error difference positively correlates with JS distance. In other words, the advantage of the prospective rehearsal becomes more pronounced the more training and test distribution deviate. This is in line with our theoretical analysis: The more training and test distribution differ, the more prediction errors of $f$ we expect, which means that our proposed prospective rehearsal helps more. Since real-world prosthesis data is prone to noise and variability from various exteroceptive and interoceptive factors (Prahm et al., 2019), being robust to distribution shifts is an important requirement for this domain, and our approach performs well in this regard.
@@
Resilience to adversarial perturbations. To further assess the resilience of the prospective rehearsal approach, we compare its outcomes to those without it in the face of adversarial perturbations. These adversarial samples are created using the Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2014) across diverse epsilon thresholds, $\tau$. Notably, our prospective rehearsal technique surpasses the conventional baseline across all evaluated FGSM attack intensities (Figure 4A). Furthermore, as the perturbation intensity escalates, the performance disparity between the prospective and the conventional method widens, reinforcing the superior adaptability of prospective rehearsal under scenarios where test data distribution deviates from the training set. The results are statistically significant.
@@
Gait stability assessment with an exoskeleton. In a pilot study, able-bodied individuals walked with an ankle exoskeleton (Figure 5 left) controlled by joint profiles predicted by the model with the prospective rehearsal approach. Participants reported that the predictions felt timely, accurate, and in sync with their residual body movements. To evaluate the stability of the model-based foot profiles, with and without the prospective rehearsal approach, we measured the maximal Lyapunov exponent using the Eckmann method (Eckmann et al., 1986). The Lyapunov exponent measures the rate of divergence of nearby trajectories in a dynamical system, with lower values indicating higher stability. Our results showed that both the first and second Lyapunov exponents were lower when using our proposed method, indicating more stable gait trajectories (Figure 5 right).

### v4

@@ 1 Introduction
Lower limb amputation and neuromuscular disorders detrimentally affect natural locomotion, compromising individuals’ quality of life [1]. Those afflicted often depend on assistive technologies like prosthetics or orthoses to regain daily mobility [1]. However, conventional passive prosthetics often fall short in replicating natural gait during diverse activities, from basic transitions like standing from a seated position to more dynamic actions like running or navigating slopes [1]. The advent of powered prosthetics, equipped with integrated motors, promises a more naturalistic gait. Nonetheless, this advancement requires a model capable of effectively approximating the complexities of human gait synergy to accurately estimate the necessary motor commands. While finite-state machines are adequate for rudimentary scenarios [2; 3], their construction becomes infeasible for finer gait phase resolution and multiple locomotion tasks [4]. An emerging solution lies in training models on human demonstrations to intuitively infer target limb motion [5; 6; 7]. This methodology not only replicates gait’s inherent fluidity but also facilitates seamless transitions across varying gaits, eliminating the need for rigid rules or heuristics. Furthermore, such learning-based gait models can be robustified by training on a large and diverse set of human demonstrations [8]. To achieve personalization, these models could incorporate references that exhibit anthropomorphic features similar to those of the physically impaired subject [9].
@@
Recent studies attest to the efficacy of learning-based models in gait and prosthetic behavior estimation [10; 11; 12; 13]. However, most of these works have focused on either a single locomotion task or jointly training multiple modes in the context of multi-locomotion scenarios. Furthermore, prior research has approached multitask learning from human demonstration as a standard supervised learning task, which assumes independence across time. This oversight can cascade predictive errors, deviating significantly from the original training distribution [14; 15]. Some approaches [16; 17; 18] have addressed the issue of state distribution shift by using a trained model to collect additional data and label them as the model encounters new states, refining the model iteratively. However, these studies require having access to a simulator, environment, or deployment of models in the real world for data collection to estimate the expert action for the induced state. This may not be feasible or highly costly for multiple safety-critical scenarios, e.g., human-centered robotics that might entail risks of imbalance and fall. Moreover, these approaches have largely focused on single-task settings rather than continual multitask adaptation. The multifaceted dynamics of human gait, which can shift based on varying terrains, speeds, fatigue levels, or specific locomotion tasks, underscore the necessity for a model that fluidly adapts to these evolving conditions.
@@ 2 Related Work
Prosthetic behavior models. Conventional prosthetic behavior models primarily rely on state machines that deterministically generate control commands based on finite states within a gait cycle. These models segment the continuous gait cycle into discrete states [2; 3; 19], which fails to accurately replicate the smooth and continuous movement patterns typical of human walking. To better model the fluidity of the gait cycle, learning-based regression methods [8; 20; 10; 12; 11; 21] have been utilized that directly predict the joint motion commands from the sensor states of residual body movement. However, these models generally adopt a straightforward reactive strategy, learning the correlation between the residual states of the body and the limb’s movements. Hence, these models often oversimplify the complex nature of human motor abilities, which are inherently adaptable to various tasks, diverse environments, and evolving gait patterns.
@@
Multitask adaptive learning. A principal challenge in multitask adaptive learning is the phenomenon of catastrophic forgetting [22; 23; 24], where newly acquired knowledge can cause the loss of previously learned information. To combat this, continual learning strategies, such as regularization-based methods, architectural strategies, and experience replay (ER)[25; 26; 27] have been developed. Regularization-based methods like Elastic Weight Consolidation (EWC) [25] and Synaptic Intelligence (SI) [26] introduce regularization terms to guide the model towards a parameter space optimized for low error across both previous and new tasks. Architectural methods such as Progressive Neural Networks (PNNs) [28] expand the model’s structure by introducing task-specific parameters to sustain performance across tasks. Rehearsal or replay methods, including Experience Replay (ER) [23; 29; 30] and Gradient Episodic Memory (GEM) [31], maintain a sample of previous tasks to be revisited during new task training, reinforcing the model’s exposure to the old data distribution. Rehearsal approaches have empirically proved to be the most effective among several approaches developed [32]. However, these techniques have a retrospective nature and are meant for treating data samples independently. In the context of bionic prosthetics, where predictive models interact with a dynamic environment, anticipating the impact of predictions on subsequent data inputs is critical.
@@
Model-based reinforcement learning (RL). Model-based RL focuses on developing a representation of the environment by gathering data from interactions driven by a specific policy. This constructed model of the environment facilitates a level of planning, enabling the estimation of potential outcomes for future states [33; 34; 35; 36; 37]. Typically, these methods rely on a reward signal and online interactions with the environment. With the advent of offline reinforcement [38; 39] and model-based imitation learning (IL) [40; 41], access to the reward signals or online interactions can be bypassed [42]. Both model-based RL and IL, by internalizing environmental dynamics, enable an agent to forecast outcomes and make informed decisions. Our approach is inspired by model-based RL; however, we use it as a rehearsal strategy within a multitask continual learning setting to mitigate forgetting, adapt to evolving tasks, and limit compounding errors, all without needing access to a reward signal or online interactions.
@@ 3 Method
Multitask learning via generic and task-specific modules. A core challenge in multitask learning is to exploit commonalities between the tasks while maintaining flexibility to account for the particularities of every single task. Our approach is to compose the gait model $f$ of two functions: a shared backbone $f^{s}$ across tasks, followed by task-specific layers $f_{t}$. The overall prediction for an input sample $\{\mathbf{x}_{t,i-T}...\mathbf{x}_{t,i}\}$ for task $t$ at time $i$ is $f(\mathbf{x}_{t,i-T}...\mathbf{x}_{t,i})=f_{t}[f^{s}(\mathbf{x}_{t,i-T}...\mathbf{x}_{t,i})]$. This duality enables the model to distill common features across all tasks and concurrently adapt to the unique aspects of individual tasks. We employ temporal convolutions [43; 12] to model the shared backbone due to their effectiveness in handling temporal sequences. To tailor the model to specific tasks, we implement task-specific layers using a two-layer feed-forward network, providing lightweight and efficient customization for each task.
@@
Continual learning via evolving architecture and rehearsal. In (task) incremental adaptation, the model is trained incrementally with one task at a time. Thus, the model does not require data from all the tasks during training. When a new task $t$ is encountered, a new task-specific prediction layer $f_{t}$ is created and both $f_{t}$ and $f^{s}$ are trained using task-specific data. However, this method is prone to catastrophic forgetting [24; 22], that is, the model forgets the previously trained tasks when it learns a new task. To deal with the forgetting problem, we include samples from prior tasks in the training (i.e., a ’rehearsal’ [44]). However, beyond prior rehearsal approaches, we do not merely copy training data from prior tasks. Instead, we augment it, incorporating prediction errors from the current model $f$ using the following scheme.
@@
Prospective model. A principal challenge in our scenario is that prediction errors at time step $k$ may negatively impact prediction performance at time step $k+1$, such that samples are not identically and independently distributed across time. In more detail, for the human demonstrations, we know that the sensor state $\mathbf{x}_{t,k}$ and desired kinematic profile $y_{t,k}$ is followed by the sensor state $\mathbf{x}_{t,k+1}$. However, if we predict a slightly different kinematic profile $\hat{y}_{t,k}=f(\mathbf{x}_{t,k-T}...\mathbf{x}_{t,k})\neq y_{t,k}$, we will also observe a different sensor state $\hat{\mathbf{x}}_{t,k+1}\neq\mathbf{x}_{t,k+1}$ in the next time step, pushing us away from the training data distribution. Such deviations can accumulate over time, which is a well-known phenomenon in imitation learning [14; 15]. In general, imitation learning is a helpful metaphor for our situation, as we also wish to mimic the demonstration of human experts in a setting where independence over time does not hold. However, past theory on imitation learning has usually assumed constant bounds on the error in every time step. Such constant bounds are not realistic in our scenario with a continuous space. Unfortunately, if we slightly relax the assumption of constant bounds to Lipschitz bounds, the deviation between desired time series and predicted time series can grow exponentially, as we show in the following, simple theorem.
@@
This observation motivates the key innovation of our work, namely a prospective rehearsal. More precisely, we train a prospective model $g$ to mimic the dynamics of the system with $g(\mathbf{x}_{k},{y}_{k})\approx\mathbf{x}_{k+1}$ and imagine the subsequent inputs resulting from the predictions of $f$. We then add data tuples $(\textbf{x}^{\prime}_{k+1},{y}_{k+1})$ with $\textbf{x}^{\prime}_{k+1}=g\big{(}\mathbf{x}_{k},f(\mathbf{x}_{k-T}...\mathbf{x}_{k})\big{)}$ to our training data. These additional training data tuples take the prediction errors of $f$ into account and thus help $f$ to counteract its own prediction errors. In our scenario, ${y}_{k+1}$ is the right target for the input $\textbf{x}^{\prime}_{k+1}$ because ${y}_{k+1}$ represents a target joint profile in a prosthetic limb. This target profile should remain the same in the same gait phase, even if $\textbf{x}^{\prime}_{k+1}$ slightly deviates from $\mathbf{x}_{k+1}$. In a general scenario, other targets might be required. Theoretically speaking, our aim is to achieve a constant bound $d\big{(}g[\mathbf{x}_{k},f(\mathbf{x}_{k-T},...,\mathbf{x}_{k})],\mathbf{x}_{k+1}\big{)}\leq\epsilon$ for small $\epsilon>0$, such that classical imitation learning theory applies [14; 15] and an exponential deviation over time is avoided. We do so by minimizing $d\big{(}g[\mathbf{x}_{k},{y}_{k}],\mathbf{x}_{k+1}\big{)}$ explicitly in Eq. equation 1.
@@ 4 Experiments
Datasets and metrics. We use three real-world human gait experiments’ datasets. The first two, ENABL3S [45] and Embry [46], are publicly available from popular gait labs. ENABL3S comprises approximately 5.2 million training samples and 1.3 million test samples collected from ten subjects performing various locomotion activities, including level-ground walking, stair ascent and descent, and ambulating on an inclined walkway. From the Embry dataset, we selected data representing nine very distinct combinations of walking speeds (ranging from 0.8 m/s to 1.2 m/s) and inclines (ranging from -10 degrees to +10 degrees). This selection resulted in approximately 512,000 training samples and 104,000 test samples from ten subjects, encompassing a diverse range of movements. The third dataset is a novel collection from our lab, focusing on gait patterns of patients (transtibial amputees), recorded with a 200 Hz infrared motion capture system using twelve cameras, and includes activities like walking, stair ascent, and descent, with around 27K training and 11K testing samples. Ethical clearance for these experiments was obtained from our institutional review board. This dataset, along with the others, includes motion-related variables such as 3D angles of body joints and segments, velocities, and accelerations. To evaluate the accuracy of the desired joint profiles, $y_{k}$, we calculate the coefficient of determination ($R^{2}$), between the ground truth trajectories and model-based joint profiles.
@@
Ablations. We assess the contribution of each component on overall performance. We explore both joint and continual learning paradigms, comparing models with shared final layers against those with task-specific final layers [47]. Further, we consider both conventional and prospective rehearsal with each of the shared and task-specific final layer variants. In joint training, the training data is either augmented with the multitask prospective rehearsal from a validation set or simply with the original samples from the validation set itself (conventional rehearsal). In all conditions, we used the same amount of training data to maintain fairness. Augmentation with prospective rehearsal improved the performance of models with both shared and task-specific final layers. Further, the results provide evidence that task-specific final layers help to account for particularities among tasks. A comparison of performance in task-incremental learning yields similar results. The results demonstrate that the choice of a combination of prospective rehearsal and an evolving architecture, where task-specific final layers are integrated with a shared backbone, yields the best performance across the board (Table 1A–B). All following experiments are performed with task-specific final layers.
@@
Resilience to distribution shift. In some cases, we do not observe a significant improvement in the performance of the prospective rehearsal over a conventional rehearsal. To explain these results, we inspect the difference in $R^{2}$ between the prospective and conventional rehearsal versus the Jenson-Shannon (JS) distance [48] between training and test distributions (Figure 3). We observe that the error difference positively correlates with JS distance. In other words, the advantage of the prospective rehearsal becomes more pronounced the more training and test distribution deviate. This is in line with our theoretical analysis: The more training and test distribution differ, the more prediction errors of $f$ we expect, which means that our proposed prospective rehearsal helps more. Since real-world prosthesis data is prone to noise and variability from various exteroceptive and interoceptive factors [49], being robust to distribution shifts is an important requirement for this domain, and our approach performs well in this regard.
@@
Resilience to adversarial perturbations. To further assess the resilience of the prospective rehearsal approach, we compare its outcomes to those without it in the face of adversarial perturbations. These adversarial samples are created using the Fast Gradient Sign Method (FGSM) [50] across diverse epsilon thresholds, $\tau$. Notably, our prospective rehearsal technique surpasses the conventional baseline across all evaluated FGSM attack intensities (Figure 4A). Furthermore, as the perturbation intensity escalates, the performance disparity between the prospective and the conventional method widens, reinforcing the superior adaptability of prospective rehearsal under scenarios where test data distribution deviates from the training set. The results are statistically significant.
@@
Gait stability assessment with an exoskeleton. In a pilot study, able-bodied individuals walked with an ankle exoskeleton (Figure 5 left) controlled by joint profiles predicted by the model with the prospective rehearsal approach. Participants reported that the predictions felt timely, accurate, and in sync with their residual body movements. To evaluate the stability of the model-based foot profiles, with and without the prospective rehearsal approach, we measured the maximal Lyapunov exponent using the Eckmann method [51]. The Lyapunov exponent measures the rate of divergence of nearby trajectories in a dynamical system, with lower values indicating higher stability. Our results showed that both the first and second Lyapunov exponents were lower when using our proposed method, indicating more stable gait trajectories (Figure 5 right).